OpenStudy (anonymous):

Find the circumference of the circle R = 2x + 1 A. 6.28x + 3.14 B. 3.142 + 6.28 C. 12.56x + 6.28

Directrix (directrix):

C = 2 pi r C = 2 pi (2x+1) C = pi (4x + 2) C = 3.14 ( 4x + 2) C = 12.56 x + 6.28 ---> C